OpenStudy (anonymous):

The density of gasoline is 0.65 g/mL. What is the mass in pounds of 45.0L (about 12 gallons) of gasoline? How would I set up this problem?? Thank you!

OpenStudy (lgbasallote):

\[\rho = \frac MV\] so \[V\rho = M\] (\(\rho\) means density ) so... \[M = 0.65 g/mL \times 45.0 L \implies \boxed{0.65 g/mL \times 45,000 mL}\] does that help?